OpenStudy (anonymous):

-(-8)-3I7-11I << supposed to be the absolute value of 7 - 11

OpenStudy (anonymous):

\[-(-8)-3|7 - 11|\]

OpenStudy (anonymous):

@lgbasallote @ganeshie8

OpenStudy (anonymous):

\[R= G+T/ G-T solve for T\]

OpenStudy (calculusfunctions):

-(-8) − 3|7 − 11| = 8 − 3|-4| = 8 − 3(4) because the absolute value of any number is always positive. = 8 −12 = -4

OpenStudy (anonymous):